# DeCo: Defect-Aware Modeling with Contrasting Matching for Optimizing Task Assignment in Online IC Testing

### v1

@@ 1 Introduction
Current studies primarily concentrate on automating fault localization within IC testing Fan et al. (2020)Fan et al. (2023), with the goal of pinpointing potential fault locations or defect types during manufacturing, or improving the detection yield during the design phase to reduce both cost and time He and Wang (2007)Lang et al. (2022). However, focusing solely on fault localization may lead to critical limitations. ❶ Specifically, certain defect types exhibit similar failure characteristics, and when defect data are scarce, identifying them becomes inherently challenging. As such, relying solely on automated fault localization is inadequate, as it does not utilize contextual insights from previous cases. Engineers with prior experience in similar failure scenarios are still often required to ensure accurate defect diagnosis. ❷ In additions, existing approaches overlook the structural relationships between defects, historical failure logs, and engineers’ workload and expertise, which hinder effective RMA handling by leading to suboptimal task assignments and inefficient defect resolution. These limitations reveal a research gap: existing work focuses on defect detection, but lacks research on optimizing RMA handling to reduce manual diagnosis time and enhance task allocation efficiency as shown in Figure 1.
Addressing this gap is essential for reducing resolution time and enhancing overall IC testing workflows.
To address the identified challenges and enhance RMA handling, we propose DeCo (Defect-Aware Modeling with Contrasting Matching), a novel framework designed to optimize RMA task assignment by maximizing task-handling success rates, while ensuring engineer workloads remain manageable. Specifically, to overcome the first limitation, we design a defect-aware graph based on ATE logs of each RMA task, structuring co-failure relationships among RMAs. This graph structure encodes hidden dependencies among defects, enabling more accurate defect differentiation even when ATE logs for defect types are scarce. Additionally, we formulate defect-aware representations for engineers and RMA tasks based on their relationships with defect types. To further reinforce representations, we design local and global structure modeling on the defect-aware graph, effectively capturing failure characteristics. These two modeling approaches can be deployed independently or integrated to provide a more comprehensive understanding of defect characteristics, directly addressing the second limitation. Finally, a contrasting-based assignment is introduced to match new RMA tasks with the most suitable engineers based on their representations and engineers’ current workloads, ensuring optimal task assignment and improving overall efficiency in RMA handling.
@@ 2.1.2 Analysis on Failure Characteristics
We perform a preliminary data analysis to validate that different defect types may share similar failure characteristics, as mentioned in Sec. 1. Figure 2 shows the fail rate of each test name across defect types in all ATE logs. We can observe that while fail rates vary, some defect types exhibit similar failure characteristics, whereas others differ significantly.
@@ 3.1.1 Engineer Representation
To evaluate the experience and ability of each engineer in handling different kinds of defect types of ICs, we first formulate the normalized representation $\bar{\textbf{e}}_{i}$ for each engineer $e_{i}\in E$, utilizing the data from RMAs they have previously handled. Given the defect type set $\mathcal{V}_{type}$, let $\mathcal{P}_{i,p}$ and $\mathcal{N}_{i,p}$ denote the average processing time and the count of defect type $p\in\mathcal{V}_{type}$ handled by engineer $e_{i}\in E$, respectively. An engineer $e_{i}$’s ability to handle defect type $p$ is defined as follows, where shorter processing time and more handling experience indicate higher ability:
${\small\text{ability}(i,p)=\text{avg}\biggl{(}1-\frac{\mathcal{P}_{i,p}}{\sum_{e_{j}\in E}\mathcal{P}_{j,p}/|E|}\>,\>\frac{\mathcal{N}_{i,p}}{\sum_{e_{j}\in E}\mathcal{N}_{j,p}/|E|}\biggr{)}.}$ (1)
@@ 3.1.2 Task Representation
Local structure modeling. For local structure modeling, we represent the characteristics of each node in the defect-aware graph $\mathcal{G}$ by exploring its local neighborhood. This allows us to encode fine-grained, proximity-based relationships among test items, tasks, and defect types. To achieve this, we redesign the classic model DeepWalk Perozzi et al. (2014), which learns node embeddings by generating random walks over the graph and optimizing node co-occurrence within local neighborhoods. Also, we extend the skip-gram architecture Grover and Leskovec (2016)Mikolov et al. (2013) by considering the second-order proximity Tang et al. (2015) of each node in defect-aware graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$ to better capture structural proximity. The objective function of local structure modeling is designed as follows:
@@
${\small\omega_{v,u}=\begin{cases}J\bigl{(}N(v)\cap\mathcal{V}_{test},N(u)\cap\mathcal{V}_{test}\bigr{)}&\text{, if }v\neq u\text{ and }v,u\in\mathcal{V}_{task}\\ 0&\text{, otherwise }\end{cases},}$ (5)
where $Z_{u}=\sum_{u\in\mathcal{V}}\exp(\psi(v)\cdot\psi(u))$, and $N(u)$ refers to th set of nodes in the random walk starting from $u$. The function $J(\cdot,\cdot)$ represents the Jaccard similarity of two sets. $\psi(v)\in\mathbb{R}^{d}$ and $\psi(u)\in\mathbb{R}^{d}$ are embedding vectors of nodes $v$ and $u$, respectively.
@@ 3.2 Contrasting-Based Assignment
After obtaining the defect-aware representation $\bar{\mathbf{e}}_{i}$ (Eq. 2) and $\bar{\mathbf{t}}_{k}$ (Eq. 10) for each engineer $e_{i}$ and newly received task $t_{k}$, we aim to assign the most appropriate engineer to handle the new task $t_{k}$. Here, we design the assignment score $\mu_{i,k}$ to quantify the suitability of each engineer $e_{i}\in E$ for new task $t_{k}$. Inspired by contrastive decoding approaches Li et al. (2022)Chuang et al. (2023), the assignment score $\mu_{i,k}$ is computed by contrasting information across representations $\bar{\mathbf{t}}_{k}$ and $\bar{\mathbf{e}}_{i}$ to refine the task-to-engineer assignment process, as designed as follows:
@@ 4.1 Experimental Setup
Dataset and Preprocessing. We use a real-world ATE dataset collected from a semiconductor company, as the format described in Sec. 2.1.1. To evaluate methods under both scarce and informative defect data conditions, we consider the original dataset as scarce and create an expanded dataset by increasing the number of engineers, defect types, RMA logs and other relevant factors. The statistics of the scarce (original) dataset and the expanded dataset are presented in Table 2. The first 75% of the total RMA data is used for training, while the remaining 25% is reserved for testing.
Table 2: Dataset statistics.
@@
Evaluation Metrics.
@@ 4.2 Results and Analysis
Comparison Results. The comparative results are presented in Table 3. To evaluate different simulations of engineer processing time, we report the results with $\kappa=\{1,2,5\}$ (Eq. 13). The key observations are as follows:
@@
Performance on Defect Type Detection. Besides task assignment, we evaluate DeCo for defect type detection using defect-aware representations. In Eq. 9, $\text{prob}(k,p)$ denotes the probability of task $t_{k}$ belonging to defect type $p$. We select the top-3 and top-5 defect types with the highest $\text{prob}(k,p)$ values as detection results. Figure 5 presents the results, analyzing the impact of varying walk length and margin in DeCo. The results show that better performance (a stronger red color) is achieved when both walk length and margin are either smallest or largest. For improvements across all metrics, increasing walk length enhances performance when margin is set to 10, and increasing margin helps only when walk length is at its maximum. This implies that while longer walks capture more informative structures, a larger margin is also required to maintain proper separation in the embedding space. Finding an appropriate balance between walk length and margin can further enhance performance.
@@
Case 1. Figure 6a shows a task which belongs to defect type “#EIPD”. Based on the learned representation $\bar{\textbf{t}}$ of the task, the defect type can be correctly detected. DeCo then assigns this task to Engineer A, whose defect-aware representation $\bar{\textbf{e}}$ indicates the highest capability for handling “#EIPD”. Engineer A also has the highest historical handling ratio for this defect type. This case demonstrates that defect-aware representations learned in DeCo can effectively capture failure characteristics, enabling accurate defect detection and task assignment to the most experienced engineers.
@@ 5 Related Work
Fault Localization. Research on IC testing primarily focuses on optimizing test design in hardware design and manufacturing processes Lang et al. (2022); Park et al. (2017); Afacan et al. (2021); Lyu et al. (2018). Specifically, Fan et al. (2020) adopted machine learning methods for equipment monitoring, fault detection, and diagnosis during the thin film deposition stage. Fan et al. (2023) used VAE-based generative models to address data imbalance in thin film deposition. Liau and Schmitt-Landsiedel (2003) focused on detecting and analyzing power supply noise in ICs using ATE.
Although the above works perform well in various scenarios, certain defects may exhibit similar failure characteristics, making identification challenging. In such cases, accurate diagnosis often relies on experienced engineers. Therefore, our work focuses on identifying IC failure characteristics and assign them to suitable engineers to improve defect handling. This is an area still unexplored in IC test research.
Task Assignment. Recent research focuses on effective assignment while considering constrains Mo et al. (2013); Hettiachchi et al. (2022); Dickerson et al. (2018); Andersen et al. (2016); Liu et al. (2015); Constantino et al. (2017). Specifically, Zheng et al. (2015) developed the QASCA system, focusing on quality-aware task assignment and considering worker skills and quality standards. Various evaluation metrics are also used depending on the task type Li et al. (2014); Lee et al. (2014); Balakrishnan et al. (2015); De Alfaro and Shavlovsky (2014). Mavridis et al. (2016) assessed the time required for workers to complete tasks, ensuring timely task completion. Goel and Faltings (2019) considered fairness among the population under budget constraints.
While these works address worker skills, task difficulty, and efficiency, they often overlook worker workloads, a critical factor in real-world applications. Therefore, in our RMA task assignment problem, we not only consider engineers’ expertise and task failure characteristics but also aim to maximize the task handling success rate while maintaining a balanced workload, ensuring practical applicability.
@@ 6 Conclusion
We propose the DeCo framework for optimizing RMA task assignment. To capture failure characteristics from ATE logs, DeCo constructs a defect-aware graph and employs two novel modeling strategies: local and global structure modeling to learn comprehensive task representations. A contrast-based assignment method refines task-to-engineer allocation. When applied to real-world IC testing data, DeCo achieves the highest success rates while ensuring engineers’ workloads remain manageable across both scarce and expanded ATE log conditions. We also demonstrate its effectiveness in defect type detection. Two case studies show that DeCo can assign potentially capable engineers to previously unfamiliar defects.

### v2

@@ 1 Introduction
Current studies concentrate on automating fault localization within IC testing Fan et al. (2020, 2023), with the goal of pinpointing potential fault locations or defect types during manufacturing, or improving the detection yield during the design phase to reduce both cost and time He and Wang (2007); Lang et al. (2022). However, focusing solely on fault localization may lead to critical limitations. ❶ Specifically, certain defect types exhibit similar failure characteristics, and when defect data are scarce, identifying them becomes inherently challenging. As such, relying solely on automated fault localization is inadequate, as it does not utilize contextual insights from previous cases. Engineers with prior experience in similar failure scenarios are still often required to ensure accurate defect diagnosis. ❷ In additions, existing approaches overlook the structural relationships between defects, historical failure logs, and engineers’ workload and expertise, which hinder effective RMA handling by leading to suboptimal task assignments and inefficient defect resolution. These limitations reveal a research gap: existing work focuses on defect detection, but lacks research on optimizing RMA handling to reduce manual diagnosis time and enhance task allocation efficiency as shown in Figure 1.
Addressing this gap is essential for reducing resolution time and enhancing overall IC testing workflows. To address the identified challenges and enhance RMA handling, we propose DeCo (Defect-Aware Modeling with Contrasting Matching), a novel framework designed to optimize RMA task assignment by maximizing task-handling success rates, while ensuring engineer workloads remain manageable. Specifically, to overcome the first limitation, we design a defect-aware graph based on ATE logs of each RMA task, structuring co-failure relationships among RMAs. This graph structure encodes hidden dependencies among defects, enabling more accurate defect differentiation even when ATE logs for defect types are scarce. Additionally, we formulate defect-aware representations for engineers and RMA tasks based on their relationships with defect types. To further reinforce representations, we design local and global structure modeling on the defect-aware graph, effectively capturing failure characteristics. These two modeling approaches can be deployed independently or integrated to provide a more comprehensive understanding of defect characteristics, directly addressing the second limitation. Finally, a contrasting-based assignment is introduced to match new RMA tasks with the most suitable engineers based on their representations and engineers’ current workloads, ensuring optimal task assignment and improving overall efficiency in RMA handling.
@@ 2.1.2 Analysis on Failure Characteristics
We perform a preliminary data analysis to validate that different defect types may share similar failure characteristics, as mentioned in Sec. 1. Figure 2 shows the fail rate of each test name across defect types in all ATE logs. While fail rates vary, some defect types exhibit similar failure characteristics, whereas others differ significantly.
@@ 3.1.1 Engineer Representation
To evaluate the ability of each engineer in handling different defect types of ICs, we first formulate the normalized representation $\bar{\textbf{e}}_{i}$ for each engineer $e_{i}\in E$, utilizing the data from RMAs they have previously handled. Given the defect type set $\mathcal{V}_{type}$, let $\mathcal{P}_{i,p}$ and $\mathcal{N}_{i,p}$ denote the average processing time and the count of defect type $p\in\mathcal{V}_{type}$ handled by engineer $e_{i}\in E$, respectively. An engineer $e_{i}$’s ability to handle defect type $p$ is defined as follows, where shorter processing time and more handling experience indicate higher ability:
${\small\text{ability}(i,p)=\text{avg}\biggl{(}1-\frac{\mathcal{P}_{i,p}}{\sum_{e_{j}}\mathcal{P}_{j,p}/|E|}\>,\>\frac{\mathcal{N}_{i,p}}{\sum_{e_{j}}\mathcal{N}_{j,p}/|E|}\biggr{)}.}$ (1)
@@ 3.1.2 Task Representation
Local structure modeling. For local structure modeling, we represent the characteristics of each node in the defect-aware graph $\mathcal{G}$ by exploring its local neighborhood. This allows us to encode fine-grained, proximity-based relationships among test items, tasks, and defect types. To achieve this, we redesign the classic model DeepWalk Perozzi et al. (2014), which learns node embeddings by generating random walks over the graph and optimizing node co-occurrence within local neighborhoods. Also, we extend the skip-gram architecture Grover and Leskovec (2016); Mikolov et al. (2013) by considering the second-order proximity Tang et al. (2015) of each node in defect-aware graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$ to better capture structural proximity. The objective function of local structure modeling is designed as follows:
@@
${\small\omega_{v,u}=\begin{cases}J\bigl{(}N(v),N(u)\bigr{)}&\text{, if }v\neq u\text{ and }v,u\in\mathcal{V}_{task}\\ 0&\text{, otherwise }\end{cases},}$ (5)
where $Z_{u}=\sum_{u\in\mathcal{V}}\exp(\psi(v)\cdot\psi(u))$, and $N(u)$ refers to th set of nodes in the random walk starting from $u$. The function $J(N(v),N(u))$ denotes the Jaccard similarity between $N(v)\cap\mathcal{V}_{test}$ and $N(u)\cap\mathcal{V}_{test}$. $\psi(v)\in\mathbb{R}^{d}$ and $\psi(u)\in\mathbb{R}^{d}$ are embedding vectors of nodes $v$ and $u$, respectively.
@@ 3.2 Contrasting-Based Assignment
After obtaining the defect-aware representation $\bar{\mathbf{e}}_{i}$ (Eq. 2) and $\bar{\mathbf{t}}_{k}$ (Eq. 10) for each engineer $e_{i}$ and newly received task $t_{k}$, we aim to assign the most appropriate engineer to handle the new task $t_{k}$. Here, we design the assignment score $\mu_{i,k}$ to quantify the suitability of each engineer $e_{i}\in E$ for new task $t_{k}$. Inspired by contrastive decoding approaches Li et al. (2022); Chuang et al. (2023), the assignment score $\mu_{i,k}$ is computed by contrasting information across representations $\bar{\mathbf{t}}_{k}$ and $\bar{\mathbf{e}}_{i}$ to refine the task-to-engineer assignment process, as designed as follows:
@@ 4.1 Experimental Setup
Dataset and Preprocessing. We use a real-world ATE dataset collected from a semiconductor company, as described in Sec. 2.1.1. To evaluate methods under both scarce and informative defect data conditions, we consider the original dataset as scarce and create an expanded dataset by increasing the number of engineers, defect types, RMA logs and other relevant factors. The statistics of the scarce (original) dataset and the expanded dataset are presented in Table 2. The first 75% of the total RMA data is used for training, while the remaining 25% is reserved for testing.
@@
Evaluation Metrics. Our evaluation relies on two metrics:
@@ 4.2 Results and Analysis
Comparison Results. Table 3 presents the comparative results under different engineer processing time settings, with $\kappa={1,2,5}$ (Eq. 13). Key observations include:
@@
Performance on Defect Type Detection. We further evaluate DeCo for defect type detection using defect-aware representations. In Eq. 9, $\text{prob}(k,p)$ denotes the probability of task $t_{k}$ belonging to defect type $p$. We select the top-3 and top-5 defect types with the highest $\text{prob}(k,p)$ values as detection results. Figure 5 presents the results, analyzing the impact of varying walk length and margin in DeCo. Results indicate that performance improves (stronger red) when both walk length and margin are either smallest or largest. Longer walks boost performance when the margin is 10, while a larger margin helps only at maximum walk length. This implies that while longer walks capture more informative structures, they require a larger margin to maintain proper separation in the embedding space. Balancing these two factors can further enhance performance.
@@
Case 1. Figure 6a shows a task of defect type “#EIPD”. Based on the learned representation $\bar{\textbf{t}}$ of the task, DeCo correctly detects the defect type and assigns the task to Engineer A, whose representation $\bar{\textbf{e}}$ indicates the highest capability for handling “#EIPD”. Engineer A also has the highest historical handling ratio for this defect type. This case demonstrates that defect-aware representations learned in DeCo can effectively capture failure characteristics, enabling accurate defect detection and task assignment to the most experienced engineers.
@@ 5 Related Work
Fault Localization. Research on IC testing primarily focuses on optimizing test design in hardware design and manufacturing processes Lang et al. (2022); Park et al. (2017); Afacan et al. (2021); Lyu et al. (2018). Machine learning methods have been applied for fault detection and diagnosis during the thin film deposition stage Fan et al. (2020). VAE-based generative models have been used to address data imbalance in thin film deposition Fan et al. (2023). Power supply noise in ICs has also been detected and analyzed using ATE systems Liau and Schmitt-Landsiedel (2003).
Although existing works perform well in various scenarios, certain defects may exhibit similar failure characteristics, making identification challenging. In such cases, accurate diagnosis often relies on experienced engineers. Therefore, our work focuses on identifying IC failure characteristics and assign them to suitable engineers to improve defect handling. This is an area still unexplored in IC test research.
Task Assignment. Recent research focuses on effective assignment while considering constrains Mo et al. (2013); Hettiachchi et al. (2022); Dickerson et al. (2018); Andersen et al. (2016); Liu et al. (2015); Constantino et al. (2017). Some research develops the QASCA system to support quality-aware task assignment by incorporating worker skills and quality standards Zheng et al. (2015). Various evaluation metrics are also used depending on the task type Li et al. (2014); Lee et al. (2014); Balakrishnan et al. (2015); De Alfaro and Shavlovsky (2014). Some studies have assessed the time required for workers to complete tasks Mavridis et al. (2016), while others have considered fairness among the population under budget constraints Goel and Faltings (2019).
While these works address worker skills, task difficulty, and efficiency, they often overlook worker workloads, a critical factor in practice. Our RMA task assignment addresses this by considering engineer expertise, task failure characteristics, and workload balance while maximizing success rate and ensuring practical applicability.
@@ 6 Conclusion
We propose DeCo, a framework for optimizing RMA task assignment by constructing a defect-aware graph from ATE logs and employing local and global structure modeling to learn task representations. A contrast-based assignment method refines task-to-engineer allocation. When applied to real-world IC testing data, DeCo achieves the highest success rates while keepping engineers’ workloads manageable across both scarce and expanded ATE log conditions. We also demonstrate its effectiveness in defect type detection. Two case studies show that DeCo can assign potentially capable engineers to previously unfamiliar defects.
Acknowledgments
This paper was supported in part by National Science and Technology Council (NSTC), R.O.C., under Contract 113-2221-E-006-203-MY2, 114-2622-8-006-002-TD1 and 113-2634-F-006-001-MBK.